OpenStudy (anonymous):

Factor this expression as a difference of squares. a^2 + 4a +4 -b^2

OpenStudy (anonymous):

(a+2)^2-b^2 (a+2-b)(a+2+b)

OpenStudy (anonymous):

brackett come to chat and i will show you the fractions

OpenStudy (anonymous):

(a+2)^2-b^2 =(a+2+b)(a+2-b) use the identity a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

think there is a mistake here \[a^2+4a+4=(a+2)^2\] giving \[(a+2)^2-b^2\] get \[(a+2+b)(a+2-b)\] yes?

OpenStudy (anonymous):

i find this to be correct sat . i can't find any mistake :)

OpenStudy (anonymous):

i'm in chat now sorry it -took a mniute

OpenStudy (anonymous):

\[a\div b\]

OpenStudy (anonymous):

\[\sum \frac{1}{2^n}\]

OpenStudy (anonymous):

\[\sum_{k=1}^\infty (\frac{3}{4})^k\]

OpenStudy (anonymous):

\[\dbinom{n}{k}=\frac{n!}{k!(n-k)!}\]

OpenStudy (anonymous):

how d'u remember so much stuff? :O too good !

OpenStudy (anonymous):

here to learn it basically

OpenStudy (anonymous):

i typed what i wrote in chat so you can see it

OpenStudy (anonymous):

\[\int_1^{\infty}\frac{1}{x^{\frac{2}{3}}}dx\]

OpenStudy (anonymous):

i'm copy pasting everything in my notepad...will learn them soon :)